I don't understand this math problem. I am supposed to solve the equation...
3/r+4 - 1/r =1/r (3 over r+4 - 1 over r equals 1 over r). I'm confused on the way to solve it because you can't multiply r by anything to get r+4 (common denominator). Thanks for your help!

Answer by stanbon(75887)

3/(r+4) - 1/r =1/r
----
3/(r+4) = 2/r
----
Cross-multiply to get:
3r = 2r+8
r = 8
=============
Cheers,
Stan H.